OpenStudy (anonymous):

a coach informed his team that the average score for players on the team was m. The coach further said that the average score for the first 15 players was 70. There were 25 players on the team, and the range of the scores was from 60 to 100, inclusive. What is the greatest possible difference between the highest and lowest possible values of m?

OpenStudy (anonymous):

Okay, so according to the average of 15 you have a total of 15*70 plus the additional 10 that can have a max of 100 and min of 60. Lets assume max first--> TotalMax=15*70 + 10*100 mMax = TotalMax / 25 Now, min->> TotalMin=15*70 + 10*60 mMin = TotalMin/25 Difference is mMax - mMin = (10*(100-60))/25= 16

OpenStudy (anonymous):

thx